OpenStudy (anonymous):

The sqaure of x equals 4 times the square of y. If x is 1 more than twice y, what is the value of x? -4 -1/2 -1/4 1/4 1/2

OpenStudy (chihiroasleaf):

\[\frac{ 1 }{ 2 }\]

OpenStudy (cwrw238):

x^2 = 4y^2 x = 2y + 1 2y = x - 1 y = (x-1)/2 substitute for y in fitsy equation: x^2 = 4[(x-1)/2]^2 x^2 = (x-1)^2 x^2 = x^2 - 2x + 1 2x = 1 x = 1/2

OpenStudy (anonymous):

thank u all!

OpenStudy (cwrw238):

yw